OpenStudy (anonymous):

*Will give medal!* The life expectancy of a typical lightbulb is normally distributed with a mean of 1,975 hours and a standard deviation of 25 hours. What is the probability that a lightbulb will last between 1,950 and 2,050 hours? 0.84 0.77 0.82 0.89

OpenStudy (anonymous):

\[\begin{align*}P(1950<X<2050)&=P\left(\frac{1950-1975}{25}<\frac{X-1950}{25}<\frac{2050-1975}{25}\right)\\ &=P(-1<Z<3) \end{align*}\]

OpenStudy (anonymous):

Are you calculation the z score? Im confused @SithsAndGiggles

OpenStudy (imstuck):

Yes she is calculating the z score. That's the only way to find the probability you need.

OpenStudy (anonymous):

Yeah, you transform for a normal distribution to the standard normal distribution with the formula \[Z=\frac{X-\text{mean}}{\text{standard deviation}}\]

OpenStudy (anonymous):

Okay so the answer is between -1 and 3?

OpenStudy (anonymous):

The answer is a probability, between 0 and 1.

OpenStudy (anonymous):

So where do I go from here?

OpenStudy (anonymous):

Refer to a z-table for the probabilities, or use a calculator to compute it for you: http://www.wolframalpha.com/input/?i=P%28-1%3Cz%3C3%29

OpenStudy (anonymous):

Sop its .84?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!